OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 8x + 10y = -34 16x – 5y = -43

OpenStudy (anonymous):

8x+10y=-34 16x-5y=-43 (1) 8x+10y=-34 (2) 16x-5y=-43 8x+10y=-34 + 32x-10y=-86 ------------------ 40x=-120 --- ----- 40 40 x = -3 8x + 10y = -34 8(-3) + 10y = -34 -24 + 10y = -34 10y = -34+24 10y = 10 --- --- 10 10 y = 1 x = -3 ; y = 1